OpenStudy (anonymous):

a recipe asks for 2 and 1/3 of a cup of flour. how many cups of flour are needed. A. 8 1/3 B. 8 1/6 C. 6 1/6 D. 5 5/6

OpenStudy (anonymous):

Are you sure you typed this right? 2 1/3 cups is just 2 1/3 cups. But that's not one of the possible answers.

OpenStudy (anonymous):

sorry how much is needed for 3 1/2 recipies

OpenStudy (anonymous):

It says cups, not quarts or pints maybe? They're saying they need 2 and 1/3 cups of flour. Then they're asking you how many cups of flour. The answer is 2 1/3 but that's not one of your possibilities. So I'm guessing that the question has a typing error?

OpenStudy (anonymous):

Okay. Now it makes sense.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

sorry for the confusion

OpenStudy (anonymous):

So you need 2 1/3 for ONE recipe. To find out how much you'll need for 2 1/2 recipes, you'll need to multiply 2 1/3 by 2 1/2. I like to change the fractions into what's called improper ones so that they aren't mixed fractions. That makes it easier to multiply them. 2 1/3 = 7/3 2 1/2 = 5/2 7/3 times 5/2 = 35/6 Are you okay with multiplying fractions? (Top times top over bottom times bottom) Now change this back into a mixed fraction. 35/6 = 5 5/6 so D.

OpenStudy (anonymous):

Thank You jam333